A firm makes car seats for an automaker in two varieties: leather and fabric. The standard labor-hours for these are 1.2...

  1. A firm makes car seats for an automaker in two varieties: leather and fabric. The standard labor-hours for these are 1.2 and 1.0 hours, respectively. The seats are made in two work centers, sewing and assembly, where the average percentage of work has been 40 percent for sewing and 60 percent for assembly. The following schedule for the next six weeks is given.

Week

1

2

3

4

5

6

Leather

15

20

18

21

15

13

Fabric

65

60

62

59

65

67

  1. Using the CPOF (Capacity Planning Using Overall Factors) approach, calculate the labor-hours needed in each of the two work centers in each week.
  2. Suppose the following information is given. Using the capacity bills approach, calculate the labor-hours needed in each of the two work centers in each week.
  3. show up the calculations

Work Center

Time (hours/unit)

Leather

Fabric

Sewing

0.5

0.4

Assembly

0.7

0.6

Total

1.2

1.0
Answer #1


Sewing labor hours= (unit time of leather * demand of week i * percentage of Sewing work)
+(unit time of fabric * demand of week i *percentage of Sewing work)


Assembly labor hours= (unit time of leather * demand of week i *percentage of Assembly work)+(unit time of fabric* demand of week i *percentage of Assembly work)

Sewing labor hours

Week 1 = (0.5*15*0.4)+(0.4*65*0.4) = 13.4
  
Week 2 = (0.5*20*0.4)+(0.4*60*0.4) = 13.6

Week 3 = (0.5*18*0.4)+(0.4*62*0.4) = 13.52

Week 4 = (0.5*21*0.4)+(0.4*59*0.4) = 13.64

Week 5 = (0.5*15*0.4)+(0.4*65*0.4) = 13.4

Week 6 = (0.5*13*0.4)+(0.4*67*0.4) = 13.32


Assembly labor hours

Week 1 = (0.7*15*0.6)+(0.6*65*0.6) = 29.7

Week 2 = (0.7*20*0.6)+(0.6*60*0.6) = 30

Week 3 = (0.7*18*0.6)+(0.6*62*0.6) = 29.88

Week 4 = (0.7*21*0.6)+(0.6*59*0.6) = 30.06

Week 5 = (0.7*15*0.6)+(0.6*65*0.6) = 29.7

Week 6 = (0.7*13*0.6)+(0.6*67*0.6) = 29.58